### v2

@@ Abstract
We discuss estimation and inference of conditional treatment effects in regression discontinuity designs with multiple scores. Aside from the commonly used local linear regression approach and a minimax-optimal estimator recently proposed by Imbens and Wager (2019), we consider two estimators based on random forests—honest regression forests and local linear forests—whose construction resembles that of standard local regressions, with theoretical validity following from results in Wager and Athey (2018) and Friedberg et al. (2020). We design a systematic Monte Carlo study with data generating processes built both from functional forms that we specify and from Wasserstein Generative Adversarial Networks that can closely mimic the observed data. We find that no single estimator dominates across all simulations: (i) local linear regressions perform well in univariate settings, but can undercover when multivariate scores are transformed into a univariate score—which is commonly done in practice—possibly due to the “zero-density” issue of the collapsed univariate score at the transformed cutoff; (ii) good performance of the minimax-optimal estimator depends on accurate estimation of a nuisance parameter and its current implementation only accepts up to two scores; (iii) forest-based estimators are not designed for estimation at boundary points and can suffer from bias in finite sample, but their flexibility in modeling multivariate scores opens the door to a wide range of empirical applications in multivariate regression discontinuity designs.
Keywords: Monte Carlo simulation; honest regression forests; local linear forests; nonparametric methods
@@ 1 Introduction
First introduced by Thistlethwaite and Campbell (1960) to study the effect of scholarship programs on subsequent academic outcomes, regression discontinuity (RD) designs are widely used to estimate causal effects in non-experimental settings. Sharp RD designs explore a particular treatment assignment mechanism in which individuals receive the treatment if an observed running variable scores above some known threshold. Under the potential outcomes framework, Hahn et al. (2001) show that the treatment is as good as randomly assigned in the population arbitrarily close to the assignment threshold if the conditional expectation functions (CEF) of the potential outcomes, with and without the treatment, are continuous at the cutoff. See, for example, Imbens and Lemieux (2008) and Cattaneo and Titiunik (2022) for early and recent reviews on RD designs.
@@
Estimation and inference of treatment effects in this context are, therefore, a local problem in the neighborhood of the assignment cutoff. Following Hahn et al. (2001) and Porter (2003), local polynomial regressions remain the most popular estimation strategy in the empirical RD literature. At the heart of these local methods is bandwidth selection for the weighting kernel. In designs with a univariate running variable, Imbens and Kalyanaraman (2012) propose a data-driven algorithm that yields an optimal bandwidth by minimizing the asymptotic mean squared error (MSE) evaluated at the score cutoff. This procedure, however, leads to a non-negligible bias term that does not vanish in the limit. Calonico et al. (2014) put forth a bias-corrected estimator that first subtracts from the point estimate an estimated bias term and then adjusts the standard error accordingly so that valid inference can be conducted.
@@
In many situations, the treatment status is determined by more than just one criterion. Yet univariate local linear regressions do not readily generalize to the case with multiple scores: kernel-based methods are less useful when the dimension exceeds two or three (see, for example, the discussion in Chapter 6 of Hastie et al., 2016). In addition, standard bias correction techniques for local linear regressions quickly become algebraically involved to derive and computationally difficult to implement in higher dimensions. Existing nonparametric estimators that work directly with multiple scores include methods proposed by Papay et al. (2011) and Choi and Lee (2018) that employ local linear regressions with multiple bandwidths selected by cross-validation. Zajonc (2012) extends the MSE-optimal bandwidth of Imbens and Kalyanaraman (2012) to the bivariate case. But, as pointed out by Imbens and Wager (2019), these methods yield oversmoothing bandwidths and asymptotically biased estimates, an issue also highlighted by Calonico et al. (2014). [^1] Choi and Lee (2018) also consider rule-of-thumb bandwidths that scale with $n^{-1/6}$, which are again oversmoothing in the sense that $nh^{5}\not\to 0$. To avoid the problem of selecting a multivariate bandwidth, Imbens and Wager (2019) propose an estimator based on numerical optimization that, given a second derivative bound $\mathcal{B}$ on the CEF of potential outcomes, is minimax-optimal among all estimators that are linear in the outcome variable over a class of CEF with second derivatives bounded by $\mathcal{B}$. This method, however, requires estimating the second derivative bound $\mathcal{B}$ and, as we show in Section 4, its performance depends crucially on accurate estimates of the nuisance parameter $\mathcal{B}$.
In this paper, we conduct a systematic Monte Carlo simulation to evaluate existing approaches to multivariate RD designs, including the minimax-optimal estimator of Imbens and Wager (2019) and the commonly used local linear regression approach with a transformed univariate score: to sidestep the complications of running local linear regressions in higher dimensions, empirical researchers often reduce the problem to a univariate one. For example, unemployment insurance programs usually set multiple eligibility criteria, such as age and prior labor force history in determining the maximum duration of benefits, but Schmieder et al. (2012) subset the data to consider only individuals with eligible employment histories within their respective age groups, and use age as the effective running variable. A similar subsetting approach has been taken by Jacob and Lefgren (2004) and Matsudaira (2008) to evaluate education policies where treatment eligibility is determined by standardized test scores on multiple subjects. Another common approach is to construct a univariate transformation of the multivariate score, especially in cases where the treatment boundary takes arbitrary shapes and finding a suitable subset of the data where a univariate score applies is infeasible. For example, to exploit geographic discontinuities, Black (1999) considers the shortest distance to the boundary of a geographic region; Keele and Titiunik (2015) transform bivariate coordinates to scalars by taking their rescaled Euclidean (chordal) distance to a given boundary point, and then run local linear regressions. Other examples include Battistin et al. (2009) and Clark and Martorell (2014) that use the minimum of multiple scores as the effective running variable in designs where the treatment status is determined by any of the scores crossing the threshold. These approaches, however, have several drawbacks. Considering only subsets of the data leaves unexplored a rich set of treatment effects that could be identified under a valid multivariate RD design, where the conditional average treatment effect as a multivariate function of any point along the treatment boundary is identified; see Section 2. Heterogeneous treatment effects of this kind can be important to answer policy questions. Besides, subsetting does not fully exploit signals from multiple dimensions and could suffer a loss of precision (Imbens and Wager, 2019).
Perhaps more precarious is the case where observations with the same distance to the cutoff of interest are collapsed to the same point on the real line: transforming a multivariate score may incur unexpected behaviors of the resulting density of the effective univariate score, violating key assumptions needed for both identification and estimation. In particular, the density of the transformed univariate score may be zero at the new cutoff. In Section 3.1.1, we formally discuss the issue of zero density and provide a characterization of bivariate densities with this problem. We also show in simulations that such problem can make local linear regressions biased and undercover the true treatment effect, even if the sample size gets large. Nevertheless, these aforementioned issues that could arise when reducing a multivariate RD design to a univariate one have not received much attention in empirical RD research and are difficult to detect in practice without knowing the true data generating process (DGP).
What should an empirical researcher do in the presence of multiple scores, then? We consider, as alternatives to these existing approaches, two machine-learning-type estimators that have gained popularity in estimating heterogeneous treatment effect under unconfoundedness, but have not been fully exploited in the empirical RD literature—honest regression forests of Wager and Athey (2018) and local linear forests of Friedberg et al. (2020). Like local linear regressions, these ensemble-tree methods construct a neighborhood near a given cutoff, weighting observations by how frequently they fall into the same leaf as the cutoff across all the trees in the forest. Regression forests return the weighted average of outcomes within that neighborhood as the estimate, whereas local linear forests use these forest weights to run a weighted linear regression to further capture any smooth signals. Thus, regression forests and local linear forests are, respectively, analogues of local constant and local linear regressions, with forest weights replacing kernel weights. Since the identified conditional treatment effect at a given cutoff—or more precisely, treatment boundary point, for multivariate designs—is simply the difference between two conditional means, we mimic the standard local linear regression estimation procedure by building two forests using treated and control observations, respectively, and obtain the estimated conditional treatment effect by subtracting the the control forest estimate from the treated forest estimate, both evaluated at the same treatment boundary point. These forest-based estimators are easy to implement, allow for valid inference, and can accommodate any fixed number of scores that does not grow with the sample size.
To evaluate these nonparametric estimators, we design a systematic Monte Carlo study with DGPs built both from functional forms that we specify, and, following the proposal of Athey et al. (2021), from Wasserstein Generative Adversarial Networks (WGANs) with a high degree of complexity that can closely mimic the observed data. We find that no single estimator dominates across all simulations: (i) local linear regressions perform well in univariate settings, but can undercover when multivariate scores are transformed into a univariate score, as is commonly done in practice, possibly due to the “zero-density” issue of the collapsed univariate score at the new cutoff; (ii) the minimax-optimal estimator has good bias-variance properties when the bound on the second derivative of the CEF is accurately estimated, but its performance is sensitive to the estimation of this nuisance parameter and its current implementation only accepts up to two scores; (iii) forest-based estimators are not designed for estimation at boundary points and can suffer from bias in finite sample—though local linear forests improve upon regression forests when the underlying signal exhibits strong linear trends—but their flexibility in modeling multivariate scores opens the door to a wide range of empirical applications in multivariate RD designs.
@@
Our contribution is twofold. First, we point out a potential issue of the common practice of implementing univariate local linear regressions when the running variable is in fact multivariate, give a formal characterization of when such problem arises, and show in simulations that it can lead to biased estimates even for very large sample sizes. Second, we bring into play two additional valid, easy-to-implement estimators that can be leveraged in multivariate RD designs but have been largely overlooked in empirical RD research, and together with a systematic Monte Carlo evaluation of existing estimators, provide empirical researchers with a menu of nonparametric methods that they can tailor to different empirical settings.
@@
The rest of the paper proceeds as follows. In Section 2, we begin with the identification of conditional treatment effects in multivariate RD designs. In Section 3, we review current estimation and inference practices using local linear regressions and the issue thereof, as well as the minimax-optimal estimator, followed by an introduction to honest regression forests and local linear forests. We present simulation results in Section 4. Section 5 concludes. Appendix A collects the proof of Proposition 1, and Appendix B provides additional information on the Monte Carlo study. All replication files can be accessed at github.com/yqi3/RDForest.
@@ 2 Setup and Identification
Suppose we have a sample of independent and identically distributed (i.i.d.) observations $\mathcal{S}_{n}=\{(Y_{i},X_{i})\}_{i=1}^{n}$, where $Y_{i}\in\mathbb{R}$ is the outcome of interest and $X_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d_{X}}$ is the $d_{X}$-dimensional running variable that determines the binary treatment status, $D_{i}\in\{0,1\}$, of the $i$-th unit according to some exogenous treatment assignment rule, $A:\mathcal{X}\to\{0,1\}$. If $d_{X}=1$, then $D_{i}=1$ if and only if $A(X_{i})=\mathds{1}\{X_{i}\geq x^{c}\}=1$, where $x^{c}$ is the cutoff as in the standard sharp RD design with a univariate score. In higher dimensions, the treatment assignment rule is more complex. For example, Keele and Titiunik (2015) study the effect of television advertising on election turnout, where the treatment status is determined by a geographic media market boundary. Following Zajonc (2012), let $\mathcal{T}:=\{x\in\mathcal{X}:A(x)=1\}$ denote the set of scores assigned to the treatment and let
@@
denote the treatment boundary, i.e., a point $x\in\mathcal{X}$ lies on the treatment boundary if every $\epsilon$-ball around it contains both treated and control scores. In univariate designs, $\mathscr{B}$ is simply a singleton containing the one-dimensional cutoff.
Let $\bigl{(}Y_{i}(1),Y_{i}(0)\bigr{)}$ denote the pair of potential outcomes with and without the binary treatment. We are interested in the following conditional treatment effect at any $x^{c}\in\mathscr{B}$:
@@
For each unit $i$, we only ever observe one of the potential outcomes. However, if we assume (i) the CEF $\mathbb{E}\left[Y_{i}(1)|X_{i}=x^{c}\right]$ and $\mathbb{E}\left[Y_{i}(0)|X_{i}=x^{c}\right]$ are continuous at all $x^{c}\in\mathscr{B}$, and (ii) $X$ has a density, $f_{X}(\cdot)$, that is bounded away from 0 in some neighborhood around $x^{c}$ for all $x^{c}\in\mathscr{B}$, then we can identify $\tau(x^{c})$ using neighboring observations around $x^{c}$:
@@
$=$ $~{}\mathbb{E}\left[Y_{i}|X_{i}=x^{c},D_{i}=1\right]-\mathbb{E}\left[Y_{i}|X_{i}=x^{c},D_{i}=0\right],$ (1)
where $\tau(x^{c})$ is identified as the difference between the conditional expectation of the observed outcome for the treated population and the control population, respectively, both evaluated at $x^{c}$. [^2] A detailed proof, following from a generalization of the argument in Hahn et al. (2001), can be found in Zajonc (2012, p.56). Thus, under a valid multivariate RD design, we can identify the conditional treatment effect as a function of scores along the treatment boundary $\mathscr{B}$. Such heterogeneity may be of interest to the researcher. Alternatively, Zajonc (2012) and Keele and Titiunik (2015) discuss a summary measure of the average treatment effect along the boundary, i.e., $\mathbb{E}[Y_{i}(1)-Y_{i}(0)|X_{i}\in\mathscr{B}]$, which can be estimated given a sample of boundary points by averaging the estimated conditional treatment effects at those sample boundary points. However, as pointed out by Narita and Yata (2023), the Lebesgue measure of the boundary $\mathscr{B}$ is typically zero, and hence they consider the Hausdorff measure instead and propose a two-stage least squares estimator for the average treatment effect along the treatment boundary, weighted by the Hausdorff measure. In what follows, we focus on estimating conditional treatment effects at a given treatment boundary point $x^{c}\in\mathscr{B}$ and leave how exactly to aggregate these heterogeneous treatment effects to the empirical researcher.

### v3

@@ Abstract
We discuss estimation and inference of conditional treatment effects in regression discontinuity (RD) designs with multiple scores. In addition to local linear regressions and the minimax-optimal estimator more recently proposed by Imbens and Wager (2019), we argue that two variants of random forests, honest regression forests and local linear forests, should be added to the toolkit of applied researchers working with multivariate RD designs; their validity follows from results in Wager and Athey (2018) and Friedberg et al. (2020). We design a systematic Monte Carlo study with data generating processes built both from functional forms that we specify and from Wasserstein Generative Adversarial Networks that closely mimic the observed data. We find no single estimator dominates across all specifications: (i) local linear regressions perform well in univariate settings, but the common practice of reducing multivariate scores to a univariate one can incur undercoverage, possibly due to vanishing density at the transformed cutoff; (ii) good performance of the minimax-optimal estimator depends on accurate estimation of a nuisance parameter and its current implementation only accepts up to two scores; (iii) forest-based estimators are not designed for estimation at boundary points and are susceptible to finite-sample bias, but their flexibility in modeling multivariate scores opens the door to a wide range of empirical applications, as illustrated by an empirical study of COVID-19 hospital funding with three eligibility criteria.
Keywords: Regression discontinuity designs; multiple running variables; treatment effect heterogeneity; random forests
@@ 1 Introduction
Regression discontinuity (RD) designs are widely used to estimate causal effects in observational settings, where individuals receive a binary treatment if an observed running variable is above some known threshold. Under the potential outcomes framework, Hahn et al. (2001) show that the treatment is as good as random in the subpopulation arbitrarily close to the threshold, if the conditional expectations of the two potential outcomes are continuous at the cutoff. [^1] See, e.g., Imbens and Lemieux (2008) and Cattaneo and Titiunik (2022), for reviews on RD designs. Estimation and inference of treatment effects in this context are, therefore, a local problem in the neighborhood of the threshold. Following Hahn et al. (2001) and Porter (2003), local polynomial regressions remain the most popular estimation strategy in the empirical RD literature. At the heart of these local methods is bandwidth selection for the weighting kernel. In designs with one running variable, Imbens and Kalyanaraman (2012) propose an optimal bandwidth from minimizing the asymptotic mean squared error (MSE) at the cutoff. This procedure, however, leads to an asymptotically non-vanishing bias. Calonico et al. (2014) put forth a bias-corrected estimator that first subtracts an estimated bias term and then adjusts the standard error accordingly.
@@
In many situations, the treatment status is determined by more than one criterion. Yet, univariate local linear regressions do not readily generalize to the multivariate case: kernel-based methods are less useful when the dimension far exceeds two (see, e.g., Chapter 6 of Hastie et al., 2009), and standard bias correction quickly becomes algebraically involved and computationally difficult. Existing methods for multiple scores include multiple bandwidths selected from cross-validation (Papay et al., 2011; Choi and Lee, 2018), and Zajonc (2012) extends the MSE-optimal bandwidth of Imbens and Kalyanaraman (2012) to the bivariate case. But, as pointed out by Imbens and Wager (2019), these methods yield oversmoothing bandwidths and asymptotically biased estimates; [^2] Choi and Lee (2018) also consider rule-of-thumb bandwidths that scale with $n^{-1/6}$, which are again oversmoothing in the sense that $nh^{5}\not\to 0$. to avoid selecting a multivariate bandwidth, following the literature on minimax linear estimation (e.g., Armstrong and Kolesár, 2018; Kolesár and Rothe, 2018), they propose an estimator that is minimax-optimal over a class of conditional expectations with second derivatives bounded by a given constant $\mathcal{B}$. This method, however, requires estimating $\mathcal{B}$ and, as we show in Section 4, its performance depends crucially on accurate estimation of this nuisance parameter.
@@
In this paper, we conduct a systematic Monte Carlo simulation to evaluate existing approaches to multivariate RD designs, including the minimax-optimal estimator of Imbens and Wager (2019) and the commonly used local linear regression approach with a transformed univariate score: to sidestep the complications of running local linear regressions in higher dimensions, empirical researchers often reduce the problem to a univariate one. Our contribution is twofold. First, we point out a potential issue of the common practice of implementing univariate local linear regressions when the running variable is in fact multivariate, give a formal characterization of when such problem arises, and show in simulations that it can produce biased estimates even for very large sample sizes. Second, we demonstrate two additional valid, easy-to-implement estimators that can be leveraged in multivariate RD designs, and together with a systematic Monte Carlo evaluation of existing estimators and an empirical application, provide researchers with a menu of nonparametric methods that they can tailor to different empirical settings.
1.1 Related Literature
Despite the prevalence of multi-score RD designs, empirical researchers often adopt a univariate approach. For example, unemployment insurance programs usually set multiple eligibility criteria, such as age and work history in determining the maximum duration of benefits, but Schmieder et al. (2012) subset the data to consider only individuals with eligible employment histories within their respective age groups, and use age as the effective running variable (see also Nekoei and Weber, 2017, for another example). A similar subsetting approach is taken by Jacob and Lefgren (2004) and Matsudaira (2008) to evaluate education policies where treatment eligibility is determined by multiple test scores. Another common approach is to transform the multivariate score into a univariate one, especially when the treatment boundary takes arbitrary shapes and finding a suitable subset where a univariate score applies is infeasible. For example, to exploit geographic discontinuities, Black (1999) considers the shortest distance to the boundary of a geographic region; Keele and Titiunik (2015) transform bivariate coordinates by taking their rescaled Euclidean (chordal) distance to a given boundary point and then run univariate local linear regressions, and Bui et al. (2014) take a similar Euclidean transformation. Other examples include Bleemer and Mehta (2022) that use the average of two scores, and Battistin et al. (2009) and Clark and Martorell (2014) that use the minimum of multiple scores.
These approaches, however, have several drawbacks. Considering only subsets of the data leaves unexplored a rich set of treatment effects that could be identified under a valid multivariate RD design; see Section 2. Heterogeneous treatment effects of this kind can be important to answer policy questions, as exemplified by our empirical application in Section 5. Besides, subsetting does not fully exploit signals from multiple dimensions and could suffer a loss of precision (Imbens and Wager, 2019). Perhaps more precarious is the case where observations with the same distance to the cutoff are collapsed to the same point on the real line: transforming a multivariate score may incur unexpected behaviors of the resulting density of the effective univariate score, violating key assumptions needed for both identification and estimation. In particular, the density of the transformed univariate score may be zero at the new cutoff. In Section 3.1.1, we formally discuss the issue of zero density and provide a characterization of bivariate densities subject to this problem. We also show in simulations that such problem can make local linear regressions biased and undercover the true treatment effect, even when the sample size gets large. Nevertheless, these aforementioned issues that could arise when reducing a multivariate RD design to a univariate one have not received much attention in empirical RD research and are difficult to detect in practice without knowing the true data generating process (DGP).
What should an empirical researcher do in the presence of multiple scores, then? We consider, as alternatives to these existing approaches, two estimators that have gained popularity in estimating heterogeneous treatment effect under unconfoundedness, but have not been fully exploited in the empirical RD literature—honest regression forests of Wager and Athey (2018) and local linear forests of Friedberg et al. (2020). Like local linear regressions, these ensemble-tree methods construct a neighborhood (i.e., leaf) for a given cutoff, weighting observations by how frequently they fall into the same leaf as the cutoff across all the trees in the forest. Regression forests return the weighted average of outcomes within that neighborhood as the estimate, whereas local linear forests use these forest weights to run a weighted linear regression to capture smooth signals. Thus, regression forests and local linear forests are, respectively, analogues of local constant and local linear regressions, with forest weights replacing kernel weights. Since the identified conditional treatment effect at a given cutoff is simply the difference between two conditional means, we mimic the standard local linear regression approach by building two forests using treated and control observations, respectively, and estimate the conditional treatment effect by subtracting the control forest estimate from the treated counterpart, both evaluated at the same cutoff. These forest-based estimators are easy to implement, allow valid inference, and can accommodate any fixed number of scores that does not grow with the sample size.
To evaluate these nonparametric estimators, we design a systematic Monte Carlo study with DGPs built both from functional forms that we specify, and, following the proposal of Athey et al. (2021), from Wasserstein Generative Adversarial Networks (WGANs) with a high degree of complexity that can closely mimic the observed data. We find that no single estimator dominates across all simulations: (i) local linear regressions perform well in univariate settings, but can undercover when multivariate scores are transformed into a univariate index—a common practice—possibly due to the “zero-density” issue detailed in Section 3.1.1; (ii) the minimax-optimal estimator has good bias-variance properties when the bound on the second derivative of the conditional expectation function (CEF) is accurately estimated, but its performance is sensitive to the estimation of this nuisance parameter and its current implementation only accepts up to two scores; (iii) forest-based estimators are not designed for estimation at boundary points and can suffer from bias in finite sample—though local linear forests improve upon regression forests when the underlying signal exhibits strong linear trends—but their flexibility in modeling multivariate scores opens the door to a wide range of empirical applications in multivariate RD designs.
@@ 1.1 Related Literature
The rest of the paper proceeds as follows. In Section 2, we begin with the identification of conditional treatment effects in multivariate RD designs. In Section 3, we review current estimation and inference practices using local linear regressions and the issue thereof, as well as the minimax-optimal estimator, followed by an introduction to honest regression forests and local linear forests. We present simulation results in Section 4 and an empirical application in Section 5. Section 6 concludes. Appendix A collects the proof of Proposition 1, and Appendix B provides additional information on the Monte Carlo study.
@@ 2 Setup and Identification
Suppose we have a sample of independent and identically distributed observations $\mathcal{S}_{n}:=\{(Y_{i},X_{i})\}_{i=1}^{n}$, where $Y_{i}\in\mathbb{R}$ is the outcome of interest and $X_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d_{X}}$ is the $d_{X}$-dimensional running variable that determines the binary treatment status, $D_{i}\in\{0,1\}$, of the $i$-th unit according to some exogenous treatment assignment rule, $A:\mathcal{X}\to\{0,1\}$. If $d_{X}=1$, then $D_{i}=1$ if and only if $A(X_{i})=\mathds{1}\{X_{i}\geq x^{c}\}=1$, where $x^{c}$ is the cutoff as in the standard sharp RD design with a univariate score. In higher dimensions, the treatment assignment rule is more complex. For example, Keele and Titiunik (2015) study the effect of television advertising on election turnout, where the treatment status is determined by a geographic media market boundary. Following Zajonc (2012), let
@@
denote the treatment boundary, i.e., a point $x\in\mathcal{X}$ lies on the treatment boundary if every $\epsilon$-ball around it contains both treated and control scores. In univariate designs, $\mathscr{B}$ is simply a singleton containing the one-dimensional cutoff. Let $\bigl{(}Y_{i}(1),Y_{i}(0)\bigr{)}$ denote the pair of potential outcomes with and without the binary treatment. We are interested in the following conditional treatment effect at any $x^{c}\in\mathscr{B}$:
@@
For each unit $i$, we only ever observe one of the potential outcomes. However, if we assume (i) the CEFs $\mathbb{E}\left[Y_{i}(1)|X_{i}=x^{c}\right]$ and $\mathbb{E}\left[Y_{i}(0)|X_{i}=x^{c}\right]$ are continuous at all $x^{c}\in\mathscr{B}$, and (ii) $X$ has a density, $f_{X}(\cdot)$, that is bounded away from 0 in some neighborhood around $x^{c}$ for all $x^{c}\in\mathscr{B}$, then we can identify $\tau(x^{c})$ using neighboring observations around $x^{c}$:
@@
$=$ $~{}\mathbb{E}\left[Y_{i}|X_{i}=x^{c},D_{i}=1\right]-\mathbb{E}\left[Y_{i}|X_{i}=x^{c},D_{i}=0\right],$
where $\tau(x^{c})$ is identified as the difference between the conditional expectation of the observed outcome for the treated and control populations, respectively, both evaluated at $x^{c}$. [^3] A detailed proof can be found in Zajonc (2012, p.56). Thus, under a valid multivariate RD design, we can identify the conditional treatment effect as a function of scores along the treatment boundary $\mathscr{B}$. Such heterogeneity may be of interest to the researcher. Alternatively, Zajonc (2012) and Keele and Titiunik (2015) discuss a summary measure of the average treatment effect along the boundary, i.e., $\mathbb{E}[Y_{i}(1)-Y_{i}(0)|X_{i}\in\mathscr{B}]$, which can be estimated given a sample of boundary points by averaging the estimated conditional treatment effects at those sample boundary points. However, as pointed out by Narita and Yata (2023), the Lebesgue measure of the boundary $\mathscr{B}$ is typically zero, and hence they consider the Hausdorff measure instead and propose a two-stage least squares (2SLS) estimator for the average treatment effect along the treatment boundary under the Hausdorff measure. In what follows, we focus on estimating conditional treatment effects at a given treatment boundary point $x^{c}\in\mathscr{B}$ and leave how exactly to aggregate these heterogeneous treatment effects to the empirical researcher.